### v1

@@ 6.1. Tangent Oseen equations
In this section we establish the well-posedness of the tangent Oseen equations following ideas from [20, §6]. Formally speaking, the tangent Oseen equations read: given an incompressible velocity field $\overline{\mathbf{u}}$, we seek a tangential velocity field $\mathbf{u}:\Gamma\rightarrow\mathrm{R}^{d+1}$ and a pressure field $\uppi:\Gamma\rightarrow\mathrm{R}$ such that
@@
System (6.3) is interesting as it can be regarded as a more physically accurate linearization of the incompressible tangent Navier–Stokes equations (NS) than the incompressible tangent Stokes equations (S) by introducing partial convection in terms of an a-priori approximation $\overline{\mathbf{u}}$ of the unknown velocity field $\mathbf{u}$. It is also worth mentioning that in the context of numerical methods for Navier–Stokes in flat domains, Oseen-type equations naturally arise as a result of performing semi-implicit in-time discretizations of the convective term $(\nabla\mathbf{u})\mathbf{u}$.
The following preliminary result shall prove essential in establishing the well-posedness of an appropriate weak formulation for problem (6.3).
@@
Let $d\geq 2$ be the intrinsic dimension of $\Gamma$, $1<p<\infty$ and let $p^{*}=\frac{p}{p-1}$ be its dual exponent. Moreover, assume $\Gamma$ is of class $C^{3}$. Then, for each $\overline{\mathbf{u}}\in\mathbf{L}^{d}_{t,\sigma}(\Gamma)$ the linear operator $K:\mathbf{W}^{1,p}_{t}(\Gamma)\rightarrow(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ given by
$\langle K(\mathbf{u}),\mathbf{v}\rangle:=\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}\cdot\mathbf{v}=-\int_{\Gamma}(\nabla_{\Gamma}\mathbf{v})\overline{\mathbf{u}}\cdot\mathbf{u},\qquad\forall\mathbf{u}\in\mathbf{W}^{1,p}_{t}(\Gamma),\,\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma),$ (6.4)
is well-defined, continuous and compact. Moreover, the mapping $\overline{\mathbf{u}}\mapsto K$ is linear and continuous from $\mathbf{L}^{d}_{t,\sigma}(\Gamma)$ to $\mathscr{L}(\mathbf{W}^{1,p}_{t}(\Gamma),(\mathbf{W}^{1,p}_{t}(\Gamma))^{\prime})$ with bound
$\|K\|_{\mathscr{L}(\mathbf{W}^{1,p}_{t}(\Gamma),(\mathbf{W}^{1,p}_{t}(\Gamma))^{\prime})}\leq C\|\overline{\mathbf{u}}\|_{0,d;\Gamma},$ (6.5)
where $C>0$ depends only on $\Gamma$ and $p$.
@@
In view of the density of $\mathbf{C}^{2}_{t}(\Gamma)$ in both $\mathbf{W}^{1,p}_{t}(\Gamma)$ and $\mathbf{W}^{1,p^{*}}_{t}(\Gamma)$ it is enough to first consider $\mathbf{u},\mathbf{v}\in\mathbf{C}^{2}_{t}$. For these pair of functions the first integral in (6.4) is clearly well-defined, and can be rewritten with the help of integration-by-parts formula [11, Lemma 3.3] (and density arguments) and the fact that $\operatorname{div}_{\Gamma}\overline{\mathbf{u}}=0$ as
$I(\mathbf{u},\mathbf{v}):=\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}\cdot\mathbf{v}=-\int_{\Gamma}(\nabla_{\Gamma}\mathbf{v})\overline{\mathbf{u}}\cdot\mathbf{u}.$
Consequently, Hölder’s inequality and 2.6 (Gagliardo–Nirenberg–Sobolev) yields
@@
$\leq C\|\mathbf{u}\|_{1,p;\Gamma}\|\overline{\mathbf{u}}\|_{0,d;\Gamma}\|\mathbf{v}\|_{1,p^{*};\Gamma}.$
The aforementioned density of $\mathbf{C}^{2}_{t}(\Gamma)$ and linearity of $I(\mathbf{u},\mathbf{v})$ in each of its arguments imply that $K$ is well-defined and satisfies
$|\langle K(\mathbf{u}),\mathbf{v}\rangle|\leq C\|\mathbf{u}\|_{1,p;\Gamma}\|\overline{\mathbf{u}}\|_{0,d;\Gamma}\|\mathbf{v}\|_{1,p^{*};\Gamma},\qquad\forall\mathbf{u}\in\mathbf{W}^{1,p}_{t}(\Gamma),\,\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma),$
that is, $K$ is continuous and
$\|K\|_{\mathscr{L}(\mathbf{W}^{1,p}_{t}(\Gamma),(\mathbf{W}^{1,p}_{t}(\Gamma))^{\prime})}\leq C\|\overline{\mathbf{u}}\|_{0,d;\Gamma},$
@@
Regarding the compactness of $K$, we first observe that in the case $p=p^{*}=d=2$ estimate (6.6) lets us write
$\|K(\mathbf{u})\|_{(\mathbf{H}^{1}_{t}(\Gamma))^{\prime}}\leq\|\overline{\mathbf{u}}\|_{0,\Gamma}\|\mathbf{u}\|_{0,4;\Gamma},$
which together with the compact embedding $\mathbf{H}^{1}(\Gamma)\hookrightarrow\mathbf{L}^{4}(\Gamma)$ imply that $K$ is indeed compact.
For the other cases, we proceed via an approximation argument. Using the density of $\mathbf{C}^{2}_{t}(\Gamma)$ in $\mathbf{L}^{d}_{t}(\Gamma)$ and 4.3 (Helmholtz decomposition), we can construct a sequence $\{\overline{\mathbf{u}}_{n}\}\subseteq\mathbf{C}^{2}_{t,\sigma}(\Gamma)$ such that $\lim_{n\to\infty}\|\overline{\mathbf{u}}_{n}-\overline{\mathbf{u}}\|_{0,d;\Gamma}=0$. For each $n\in\mathrm{N}$, we define $K_{n}:\mathbf{W}^{1,p}_{t}(\Gamma)\rightarrow(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ as the linear and continuous operator resulting from choosing $\overline{\mathbf{u}}_{n}$ instead of $\overline{\mathbf{u}}$ in (6.4). We have that
$\|K_{n}-K\|_{\mathscr{L}(\mathbf{W}^{1,p}_{t}(\Gamma),(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime})}\leq C\|\overline{\mathbf{u}}-\overline{\mathbf{u}}_{n}\|_{0,d;\Gamma}\xrightarrow{n\to\infty}0,$
that is, $K_{n}\to K$ in $\mathscr{L}(\mathbf{W}^{1,p}_{t}(\Gamma),(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime})$. Given the higher integrability of $\overline{\mathbf{u}}_{n}$ we can easily rewrite estimate (6.6), thus obtaining
@@
where $\varepsilon>0$ is fixed but arbitrary, and $\eta_{1},\eta_{2}>0$ depend only on $\varepsilon$. In this way, using 2.8 (compact Sobolev embedding) and [12, Theorem 6.4], we deduce that each $K_{n}$ is compact. Since the space of compact operators is closed [12, Theorem 6.1] and $K_{n}\to K$ in $\mathscr{L}(\mathbf{W}^{1,p}_{t}(\Gamma),(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime})$, we conclude that $K$ is compact. ∎
@@
Let $d\geq 2$ be the intrinsic dimension of $\Gamma$, $1<p<\infty$ and let $p^{*}=\frac{p}{p-1}$ be its dual exponent. Moreover, assume $\Gamma$ is of class $C^{3}$. Then, for each $\overline{\mathbf{u}}\in\mathbf{L}^{d}_{t,\sigma}(\Gamma)$ and $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ there exists a unique pair $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ such that
@@
Moreover, there exists $C_{\overline{\mathbf{u}}}>0$, depending only on $\overline{\mathbf{u}}$, $p$ and $\Gamma$, such that
@@
$\|\mathbf{u}\|_{1,\Gamma}$ $\leq C_{1}\|\mathbf{f}\|_{(\mathbf{H}^{1}_{t}(\Gamma))^{\prime}},$ (6.9)
$\|\uppi\|_{0,\Gamma}$ $\leq C_{2}\|\mathbf{f}\|_{(\mathbf{H}^{1}_{t}(\Gamma))^{\prime}},$
where $C_{1}>0$ depends only on $\Gamma$ and $C_{2}>0$ depends only on $\Gamma$ and affinely on $\|\overline{\mathbf{u}}\|_{0,d;\Gamma}$.
@@
We first establish the well-posedness of the following reduced problem: find $\mathbf{u}\in\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)$ such that
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}+\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}\cdot\mathbf{v}=\langle\mathbf{f},\mathbf{v}\rangle,\qquad\forall\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma).$ (6.10)
We observe that upon defining the linear and bounded operators $L,\widetilde{K}:\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)\rightarrow(\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma))^{\prime}$ (cf. (6.2)) by
$\langle L(\mathbf{u}),\mathbf{v}\rangle:=\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v},\qquad\qquad\langle\widetilde{K}(\mathbf{u}),\mathbf{v}\rangle:=\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}\cdot\mathbf{v},$
@@
$(L+\widetilde{K})\mathbf{u}=\mathbf{f}.$ (6.11)
Recalling 4.5 and 4.6 about the injectivity and surjectivity of the associated adjoint operator of $L$, a direct application of A.2 (Banach–Nečas–Babuška theorem) yields the invertibility of $L$. In turn, noticing that $\widetilde{K}=\iota^{*}_{p^{*}}\circ K\circ\iota_{p}$, where $\iota_{p}:\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)\to\mathbf{W}^{1,p}_{t}(\Gamma)$ and $\iota^{*}_{p^{*}}:(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}\to(\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma))^{\prime}$ are the canonical injections, we deduce from 6.1 (linearized convective operator) that $\widetilde{K}$ is compact.
Since $L$ is bijective and $\widetilde{K}$ is compact, we deduce from the Fredholm’s alternative for reflexive Banach spaces [12, Theorem 6.6] (see also [31, Teorema 6.9]) that $L+\widetilde{K}$ has closed range and $\dim\operatorname{ker}(L+\widetilde{K})=\dim\operatorname{ker}((L+\widetilde{K})^{\mathrm{t}})<\infty$ (cf. (A.1) in Appendix A). We also know that it suffices to prove that $\ker(L+\widetilde{K})=\{\mathbf{0}\}$ or $\ker(L+\widetilde{K})^{\mathrm{t}}=\{\mathbf{0}\}$ in order to conclude that $L+\widetilde{K}$ is invertible.
@@
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}+\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}\cdot\mathbf{v}=0,\qquad\forall\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma).$
Choosing $\mathbf{v}=\mathbf{u}\in\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)\subseteq\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma)$ and using that $\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}\cdot\mathbf{u}=0$, we obtain $\|\nabla_{\Gamma}\mathbf{u}\|_{0,\Gamma}=0$. Thus, by 3.8 (Poincaré inequality), $\mathbf{u}=\mathbf{0}$.
@@
$\langle(L+\widetilde{K})^{\mathrm{t}}\mathbf{v},\mathbf{u}\rangle=\langle(L+\widetilde{K})\mathbf{u},\mathbf{v}\rangle=\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}+\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}\cdot\mathbf{v}=0,\qquad\forall\mathbf{u}\in\mathbf{W}^{1,p}_{t,\sigma}(\Gamma).$
Since $p^{*}>2>p$, we choose $\mathbf{u}=\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma)\subseteq\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)$ to arrive at $\|\nabla_{\Gamma}\mathbf{v}\|_{0,\Gamma}=0$, and so $\mathbf{v}=\mathbf{0}$.

### v2

@@ 5.1. Tangent Oseen equations
In this section, we establish the well-posedness of the tangent Oseen equations following ideas from [19, §6]. Formally speaking, the tangent Oseen equations read: given an incompressible velocity field $\overline{\mathbf{u}}$, we seek a tangential velocity field $\mathbf{u}:\Gamma\rightarrow\mathrm{R}^{d+1}$ and a pressure field $\uppi:\Gamma\rightarrow\mathrm{R}$ such that
@@
System (5.4) is interesting as it can be regarded as a more physically accurate linearization of the incompressible tangent Navier–Stokes equations (NS) than the incompressible tangent Stokes equations (S) by introducing partial convection in terms of an a-priori approximation $\overline{\mathbf{u}}$ of the unknown velocity field $\mathbf{u}$. It is also worth mentioning that in the context of numerical methods for Navier–Stokes in flat domains, Oseen-type equations naturally arise as a result of performing semi-implicit discretizations in time of the convective term $(\nabla\mathbf{u})\mathbf{u}$.
As in flat domains, the analysis of the tangent Oseen equations hinges on the “skew-symmetry” property enjoyed by the convective term $(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}$ provided $\overline{\mathbf{u}}$ is $\operatorname{div}_{\Gamma}$-free. We now make this statement more precise. Suppose $\Gamma$ is of class $C^{2}$, let $r\in(1,\infty)$ and $\overline{\mathbf{u}}\in\mathbf{L}^{r}_{t,\sigma}(\Gamma)$. Then, for each $\mathbf{u},\mathbf{v}\in\mathbf{C}^{1}_{t}(\Gamma)$, since $\operatorname{div}_{\Gamma}\overline{\mathbf{u}}=0$, definition (4.3) with $\mathbf{v}$ replaced by $\overline{\mathbf{u}}$ and $q$ replaced by $\mathbf{u}\cdot\mathbf{v}$ yields
$0=\int_{\Gamma}\overline{\mathbf{u}}\cdot\nabla_{\Gamma}(\mathbf{u}\cdot\mathbf{v})=\int_{\Gamma}\overline{\mathbf{u}}\cdot((\nabla_{\Gamma}\mathbf{u})^{T}\mathbf{v}+(\nabla_{\Gamma}\mathbf{v})^{T}\mathbf{u}),$ (5.5)
and therefore
$\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}\cdot\mathbf{v}=-\int_{\Gamma}(\nabla_{\Gamma}\mathbf{v})\overline{\mathbf{u}}\cdot\mathbf{u},\qquad\mathbf{u},\mathbf{v}\in\mathbf{C}^{1}_{t}(\Gamma).$ (5.6)
That is, by fixing $\overline{\mathbf{u}}$, the bilinear form $(\mathbf{u},\mathbf{v})\mapsto\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}\cdot\mathbf{v}$ is skew-symmetric. In general, identity (5.6) can be extended to $\mathbf{u}$ and $\mathbf{v}$ lying in appropriate Sobolev spaces such that $\mathbf{u}\cdot\mathbf{v}\in\mathrm{W}^{1,r^{*}}(\Gamma)$. The following preliminary result makes this last observation rigorous and shall prove essential in establishing the well-posedness of an appropriate weak formulation for problem (5.4).
@@
Let $p\in(1,\infty)$ and assume $\Gamma$ is of class $C^{2}$. For each $\varepsilon>0$ define
$r=r(\varepsilon,p,d):=\begin{cases}2+\varepsilon,&\text{if $p=d=2$},\\ d,&\text{otherwise}.\end{cases}$ (5.7)
Then, for each $\overline{\mathbf{u}}\in\mathbf{L}^{r}_{t,\sigma}(\Gamma)$ the linear operator $K_{\overline{\mathbf{u}}}:\mathbf{W}^{1,p}_{t}(\Gamma)\rightarrow(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ given by
$\langle K_{\overline{\mathbf{u}}}(\mathbf{u}),\mathbf{v}\rangle:=\begin{dcases}\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}\cdot\mathbf{v},&\text{if $p^{*}<d$ or $p=d=2$},\\ -\int_{\Gamma}(\nabla_{\Gamma}\mathbf{v})\overline{\mathbf{u}}\cdot\mathbf{u},&\text{if $p<d$ or $p=d=2$},\end{dcases}$ (5.8)
is well-defined, continuous and compact. Moreover, the mapping $\overline{\mathbf{u}}\mapsto K_{\overline{\mathbf{u}}}$ is linear and continuous from $\mathbf{L}^{r}_{t,\sigma}(\Gamma)$ to $\mathscr{L}(\mathbf{W}^{1,p}_{t}(\Gamma),(\mathbf{W}^{1,p}_{t}(\Gamma))^{\prime})$ with bound
$\|K_{\overline{\mathbf{u}}}\|_{\mathscr{L}(\mathbf{W}^{1,p}_{t}(\Gamma),(\mathbf{W}^{1,p}_{t}(\Gamma))^{\prime})}\leq C\|\overline{\mathbf{u}}\|_{0,r;\Gamma},$ (5.9)
where $C>0$ depends only on $\Gamma$, $p$ and, if $p=d=2$, also on $\varepsilon$.
@@
In view of the density of $\mathbf{C}^{1}_{t}(\Gamma)$ in both $\mathbf{W}^{1,p}_{t}(\Gamma)$ and $\mathbf{W}^{1,p^{*}}_{t}(\Gamma)$ (cf. 2.7) it is enough to first consider $\mathbf{u},\mathbf{v}\in\mathbf{C}^{1}_{t}(\Gamma)$. For this pair of functions the integrals in (5.8) are clearly well-defined and, on account of (5.6), coincide. Consequently, Hölder’s inequality and 2.8 (Gagliardo–Nirenberg–Sobolev) yields
@@
$\leq C\|\mathbf{u}\|_{1,p;\Gamma}\|\overline{\mathbf{u}}\|_{0,r;\Gamma}\|\mathbf{v}\|_{1,p^{*};\Gamma},$
where $q=q(\varepsilon)\in(1,\infty)$ is such that $\frac{1}{2}+\frac{1}{2+\varepsilon}+\frac{1}{q}=1$. Notice that if both conditions $p<d$ and $p^{*}<d$ are met, then Hölder’s inequality and Sobolev embeddings $\mathbf{W}^{1,p}(\Gamma)\hookrightarrow\mathbf{L}^{\frac{dp}{d-p}}(\Gamma)$ and $\mathbf{W}^{1,p^{*}}(\Gamma)\hookrightarrow\mathbf{L}^{\frac{dp^{*}}{d-p^{*}}}(\Gamma)$ imply that $\mathbf{u}\cdot\mathbf{v}\in\mathrm{W}^{1,d^{*}}(\Gamma)$ and so the aforementioned density of $\mathbf{C}^{1}_{t}(\Gamma)$ lets us extend the validity of identity (5.6) for $r=d\geq 3$ to $\mathbf{u}\in\mathbf{W}^{1,p}_{t}(\Gamma)$ and $\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma)$. Similarly, if $p=d=2$, we can extend identity (5.6) for $r=2+\varepsilon$ to $\mathbf{u},\mathbf{v}\in\mathbf{H}^{1}_{t}(\Gamma)$ upon utilizing the Sobolev embedding $\mathbf{H}^{1}(\Gamma)\hookrightarrow\mathbf{L}^{q}(\Gamma)$ for each $q\in(1,\infty)$.
This fact, in conjunction with the aforementioned density result, the linearity of $\langle K_{\overline{\mathbf{u}}}(\mathbf{u}),\mathbf{v}\rangle$ in each of its arguments and estimate (5.10) imply that $K_{\overline{\mathbf{u}}}$ is well-defined and satisfies
$|\langle K_{\overline{\mathbf{u}}}(\mathbf{u}),\mathbf{v}\rangle|\leq C\|\mathbf{u}\|_{1,p;\Gamma}\|\overline{\mathbf{u}}\|_{0,r;\Gamma}\|\mathbf{v}\|_{1,p^{*};\Gamma},\qquad\forall\mathbf{u}\in\mathbf{W}^{1,p}_{t}(\Gamma),\,\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma),$
that is, $K_{\overline{\mathbf{u}}}$ is continuous and
$\|K_{\overline{\mathbf{u}}}\|_{\mathscr{L}(\mathbf{W}^{1,p}_{t}(\Gamma),(\mathbf{W}^{1,p}_{t}(\Gamma))^{\prime})}\leq C\|\overline{\mathbf{u}}\|_{0,r;\Gamma},$
@@
Regarding the compactness of $K_{\overline{\mathbf{u}}}$, we first observe that in the case $p=p^{*}=d=2$, estimate (5.10) let us write
$\|K_{\overline{\mathbf{u}}}(\mathbf{u})\|_{(\mathbf{H}^{1}_{t}(\Gamma))^{\prime}}\leq\|\overline{\mathbf{u}}\|_{0,2+\varepsilon;\Gamma}\|\mathbf{u}\|_{0,q(\varepsilon);\Gamma}.$
This, together with the compact embedding $\mathbf{H}^{1}(\Gamma)\hookrightarrow\mathbf{L}^{q(\varepsilon)}(\Gamma)$, implies that $K_{\overline{\mathbf{u}}}$ is indeed compact.
For the other cases, we proceed via an approximation argument. Using the density of $\mathbf{C}^{1}_{t}(\Gamma)$ in $\mathbf{L}^{d}_{t}(\Gamma)$ we first construct a sequence $\{\boldsymbol{\varphi}_{n}\}\subseteq\mathbf{C}^{1}_{t}(\Gamma)$ such that $\lim_{n\to\infty}\|\boldsymbol{\varphi}_{n}-\overline{\mathbf{u}}\|_{0,d;\Gamma}=0$. Hence, by 4.3 (Helmholtz–Weyl decomposition on $\Gamma$) we have that $\overline{\mathbf{u}}_{n}:=P_{\sigma}\boldsymbol{\varphi}_{n}=\boldsymbol{\varphi}_{n}-\nabla_{\Gamma}\Delta_{\Gamma}^{-1}\operatorname{div}_{\Gamma}\boldsymbol{\varphi}_{n}\in\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)$ for each $p\in(1,\infty)$ and
$\|\overline{\mathbf{u}}_{n}-\overline{\mathbf{u}}\|_{0,d;\Gamma}=\|P_{\sigma}\boldsymbol{\varphi}_{n}-\overline{\mathbf{u}}\|_{0,d;\Gamma}=\|P_{\sigma}(\boldsymbol{\varphi}_{n}-\overline{\mathbf{u}})\|_{0,d;\Gamma}\lesssim\|\boldsymbol{\varphi}_{n}-\overline{\mathbf{u}}\|_{0,d;\Gamma}\xrightarrow{n\to\infty}0.$
Consequently, for each $n\in\mathrm{N}$, we have that
$\|K_{\overline{\mathbf{u}}_{n}}-K_{\overline{\mathbf{u}}}\|_{\mathscr{L}(\mathbf{W}^{1,p}_{t}(\Gamma),(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime})}\leq C\|\overline{\mathbf{u}}_{n}-\overline{\mathbf{u}}\|_{0,d;\Gamma}\xrightarrow{n\to\infty}0,$
that is, $K_{\overline{\mathbf{u}}_{n}}\to K_{\overline{\mathbf{u}}}$ in $\mathscr{L}(\mathbf{W}^{1,p}_{t}(\Gamma),(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime})$. Given the higher integrability of $\overline{\mathbf{u}}_{n}$ we can easily rewrite estimate (5.10), thus obtaining
@@
where $\alpha>0$ is fixed but arbitrary, and $\eta_{1},\eta_{2}>0$ depend only on $\alpha$. In this way, using 2.10 (compact Sobolev embedding) and [12, Theorem 6.4], we deduce that each $K_{\overline{\mathbf{u}}_{n}}$ is compact. Since the space of compact operators is closed [12, Theorem 6.1] and $K_{\overline{\mathbf{u}}_{n}}\to K_{\overline{\mathbf{u}}}$ in $\mathscr{L}(\mathbf{W}^{1,p}_{t}(\Gamma),(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime})$, we conclude that $K_{\overline{\mathbf{u}}}$ is compact. ∎
@@
Let $p\in(1,\infty)$ and assume $\Gamma$ is of class $C^{2}$. Let $r=r(\varepsilon,p,d)$ be as in (5.7). Then, for each $\overline{\mathbf{u}}\in\mathbf{L}^{r}_{t,\sigma}(\Gamma)$ and $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}$ there exists a unique pair $(\mathbf{u},\uppi)\in\mathbf{W}^{1,p}_{t}(\Gamma)\times\mathrm{L}^{p}_{\#}(\Gamma)$ such that
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}+\langle K_{\overline{\mathbf{u}}}(\mathbf{u}),\mathbf{v}\rangle-\int_{\Gamma}\uppi\operatorname{div}_{\Gamma}\mathbf{v}$ $=\langle\mathbf{f},\mathbf{v}\rangle,\quad$ $\forall\,\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t}(\Gamma),$
@@
Moreover, there exists $C_{\overline{\mathbf{u}}}>0$, depending only on $\overline{\mathbf{u}}$, $p$, $\Gamma$ and, if $p=d=2$, also on $\varepsilon$, such that
@@
$\|\mathbf{u}\|_{1,\Gamma}\leq C_{1}\|\mathbf{f}\|_{(\mathbf{H}^{1}_{t}(\Gamma))^{\prime}},\qquad\|\uppi\|_{0,\Gamma}\leq C_{2}\|\mathbf{f}\|_{(\mathbf{H}^{1}_{t}(\Gamma))^{\prime}},$ (5.12)
where $C_{1}>0$ depends only on $\Gamma$ and $C_{2}>0$ depends only on $\Gamma$, affinely on $\|\overline{\mathbf{u}}\|_{0,r;\Gamma}$ and, if $p=d=2$, also on $\varepsilon$.
@@
We first establish the well-posedness of the following reduced problem: given $\mathbf{f}\in(\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma))^{\prime}$, find $\mathbf{u}\in\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)$ such that
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}+\langle K_{\overline{\mathbf{u}}}(\mathbf{u}),\mathbf{v}\rangle=\langle\mathbf{f},\mathbf{v}\rangle,\qquad\forall\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma).$ (5.13)
We observe that upon defining the linear and bounded operators $L,\widetilde{K}:\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)\rightarrow(\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma))^{\prime}$ by
[…]
@@
$(L+\widetilde{K})\mathbf{u}=\mathbf{f}.$
Applying 4.5 (injectivity of the adjoint operator) and 4.6 (surjectivity of the adjoint operator) to $L$, along with [5, Theorem A.2] (Banach–Nečas–Babuška theorem) yields the invertibility of $L$. In turn, noticing that $\widetilde{K}=\iota^{*}_{p^{*}}\circ K_{\overline{\mathbf{u}}}\circ\iota_{p}$, where $\iota_{p}:\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)\to\mathbf{W}^{1,p}_{t}(\Gamma)$ and $\iota^{*}_{p^{*}}:(\mathbf{W}^{1,p^{*}}_{t}(\Gamma))^{\prime}\to(\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma))^{\prime}$ are the canonical injections, we deduce from 5.1 (linearized convective operator) that $\widetilde{K}$ is compact.
Since $L$ is bijective and $\widetilde{K}$ is compact, we deduce from the Fredholm’s alternative for reflexive Banach spaces [12, Theorem 6.6] (see also [30, Teorema 6.9]) that $L+\widetilde{K}$ has closed range and $\dim\operatorname{ker}(L+\widetilde{K})=\dim\operatorname{ker}((L+\widetilde{K})^{\mathrm{t}})<\infty$ (cf. (A.1) in Appendix A). We also know that it suffices to prove that $\ker(L+\widetilde{K})=\{\mathbf{0}\}$ or $\ker(L+\widetilde{K})^{\mathrm{t}}=\{\mathbf{0}\}$ in order to conclude that $L+\widetilde{K}$ is invertible.
@@
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}+\langle K_{\overline{\mathbf{u}}}(\mathbf{u}),\mathbf{v}\rangle=0,\qquad\forall\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma).$
Choosing $\mathbf{v}=\mathbf{u}\in\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)\subseteq\mathbf{H}^{1}_{t,\sigma}(\Gamma)\subseteq\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma)$ and using that $\langle K_{\overline{\mathbf{u}}}(\mathbf{u}),\mathbf{u}\rangle=\int_{\Gamma}(\nabla_{\Gamma}\mathbf{u})\overline{\mathbf{u}}\cdot\mathbf{u}=0$, we obtain $\|\nabla_{\Gamma}\mathbf{u}\|_{0,\Gamma}=0$. Thus, by 3.3 (Poincaré inequality), $\mathbf{u}=\mathbf{0}$.
@@
$\langle(L+\widetilde{K})^{\mathrm{t}}\mathbf{v},\mathbf{u}\rangle=\langle(L+\widetilde{K})\mathbf{u},\mathbf{v}\rangle=\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}+\langle K_{\overline{\mathbf{u}}}(\mathbf{u}),\mathbf{v}\rangle=0,\qquad\forall\mathbf{u}\in\mathbf{W}^{1,p}_{t,\sigma}(\Gamma).$
Since $p^{*}>2>p$, we choose $\mathbf{u}=\mathbf{v}\in\mathbf{W}^{1,p^{*}}_{t,\sigma}(\Gamma)\subseteq\mathbf{H}^{1}_{t,\sigma}(\Gamma)\subseteq\mathbf{W}^{1,p}_{t,\sigma}(\Gamma)$ to arrive at $\|\nabla_{\Gamma}\mathbf{v}\|_{0,\Gamma}=0$, and so $\mathbf{v}=\mathbf{0}$.